### v2

# MP-HSIR: A Multi-Prompt Framework for Universal Hyperspectral Image Restoration

Abstract
Hyperspectral images (HSIs) often suffer from diverse and unknown degradations during imaging, leading to severe spectral and spatial distortions. Existing HSI restoration methods typically rely on specific degradation assumptions, limiting their effectiveness in complex scenarios. In this paper, we propose MP-HSIR, a novel multi-prompt framework that effectively integrates spectral, textual, and visual prompts to achieve universal HSI restoration across diverse degradation types and intensities. Specifically, we develop a prompt-guided spatial-spectral transformer, which incorporates spatial self-attention and a prompt-guided dual-branch spectral self-attention. Since degradations affect spectral features differently, we introduce spectral prompts in the local spectral branch to provide universal low-rank spectral patterns as prior knowledge for enhancing spectral reconstruction. Furthermore, the text-visual synergistic prompt fuses high-level semantic representations with fine-grained visual features to encode degradation information, thereby guiding the restoration process. Extensive experiments on 9 HSI restoration tasks, including all-in-one scenarios, generalization tests, and real-world cases, demonstrate that MP-HSIR not only consistently outperforms existing all-in-one methods but also surpasses state-of-the-art task-specific approaches across multiple tasks. The code and models are available at https://github.com/ZhehuiWu/MP-HSIR.
1 Introduction
Figure 1: PSNR comparison with the state-of-the-art all-in-one and task-specific methods across 9 tasks. Average results across all test datasets. Best viewed in color.
Hyperspectral images (HSIs) provide significantly higher spectral resolution than RGB images, as illustrated in Figure 10 (a), making them essential for urban planning [22, 62], agricultural production [56, 42], and environmental monitoring [2, 59]. However, due to various environmental factors and imaging limitations, HSIs are often affected by degradations during acquisition, compromising image quality and limiting subsequent analysis. Consequently, HSI restoration is a critical step in the image processing pipeline.
Despite advancements in task-specific HSI restoration methods [21, 78, 63, 69, 9, 35], the diversity and complexity of degradations pose a significant challenge for a single model to handle multiple HSI restoration tasks. As illustrated in Figure 10 (b), different degradation types have distinct effects on the spectral features of HSIs. For instance, noise increases spectral fluctuations, compression reduces spectral reflectance in certain bands, and haze shifts the spectral curve globally. These diverse impacts hinder the generalization capability of task-specific methods across different scenarios. Therefore, a universal framework is required to address the unique distortions caused by different degradations.
Figure 2: Overview of HSI characteristics and prompt-based all-in-one restoration methods. (a) Differences between HSIs and RGB images. (b) Impacts of different degradations on spectral features. (c) Visual prompt-based models. (d) Textual prompt-based models. (e) Dual-modality prompt-based models. (f) The proposed model integrating spectral, textual, and visual prompts.
In recent years, all-in-one image restoration methods have achieved significant progress in RGB images [32, 26, 65]. Among these, prompt learning-based approaches [52, 36, 12, 19, 1, 44] have garnered attention for their flexibility in task adaptation and efficient degradation perception. Based on prompt modality, these methods can be categorized into visual, textual, and dual prompts. As shown in Figure 10 (c), visual prompt-based models [52, 36] guide restoration by introducing additional prompt components but suffer from limited interpretability. Figure 10 (d) illustrates textual prompt-based models [12, 19], which convert textual prompts into latent embeddings to assist restoration. Nevertheless, inconsistencies may arise when the semantic gap between text and images is too large. Figure 10 (e) depicts dual prompt-based models [1, 44], which leverage pre-trained vision-language models (VLMs) to extract deep representations from text instructions and degraded images, providing more precise degradation guidance. However, existing VLMs still exhibit limitations in modeling HSI degradation information. Specifically, most all-in-one models fail to consider the unique spectral characteristics of HSIs, often resulting in spectral distortions in the restoration process.
To address the challenges of all-in-one HSI restoration, we propose a universal multi-prompt framework, MP-HSIR, which integrates spectral, textual, and visual prompts, as illustrated in Figure 10 (f). In this framework, spectral prompts guide the spatial-spectral transformer for spectral recovery, while textual and visual prompts synergistically generate degradation-specific information to guide the global restoration. Specifically, the prompt-guided spatial-spectral transformer comprises two core components: spatial self-attention and dual-branch spectral self-attention. The spatial self-attention captures non-local spatial similarities, while the global spectral branch models long-range dependencies in the spectral domain. Meanwhile, the local spectral branch incorporates universal low-rank spectral patterns from the spectral prompts to enhance spectral feature reconstruction. Additionally, the text-visual synergistic prompt deeply fuses the semantic representations of textual prompts with the fine-grained features of visual prompts, enabling controlled restoration while embedding degradation information at the pixel level. As shown in Figure 8, our method outperforms existing all-in-one methods across 9 HSI restoration tasks and surpasses state-of-the-art task-specific methods in multiple tasks. Overall, our main contributions are as follows:
  • $\bullet$
    We propose a novel multi-prompt framework for all-in-one HSI restoration, integrating spectral, textual, and visual prompts. The spectral prompt provides universal low-rank spectral patterns to enhance spectral reconstruction.
  • $\bullet$
    We introduce a text-visual synergistic prompt that combines semantic representations from textual prompts with fine-grained features from visual prompts, enhancing the controllability, interpretability, and degradation adaptability of the HSI restoration process.
  • $\bullet$
    Extensive experiments on 9 HSI restoration tasks and real-world scenarios demonstrate that MP-HSIR significantly outperforms compared all-in-one methods and surpasses state-of-the-art task-specific approaches in multiple tasks.
2 Related Work
2.1 Hyperspectral Image Restoration
Traditional HSI restoration methods mainly leverage prior knowledge, such as non-local similarity [46], low-rank structure [21, 39], and sparsity [70, 68], to constrain the solution space and employ optimization algorithms [57, 10] to solve the inverse problem. Deep learning-based HSI restoration methods directly learn the mapping between degraded and original images through deep neural networks, including convolutional neural networks (CNNs) [75, 63], transformers [33, 34, 74], and generative models [35]. Although these methods have demonstrated strong restoration capabilities, they are typically designed for specific types of degradation, such as noise removal [29, 67, 6], super-resolution [78, 64], and dehazing [45, 69], lacking universality and adaptability in practical applications.
Recent studies [74, 48, 28, 50] have explored using a single model for multiple HSI restoration tasks. DDS2M [48] designs a self-supervised diffusion model, while HIR-Diff [50] projects HSI into RGB space to leverage a pre-trained diffusion model for restoration. Nevertheless, these approaches frequently demand intricate hyperparameter tuning and exhibit slow inference speeds. Furthermore, they fail to adequately account for the distinct characteristics of different degradation types, limiting their effectiveness.
2.2 All-in-One Image Restoration
All-in-one image restoration [26, 65, 44] aims to establish a unified deep learning framework for restoring images with various degradations. However, developing an effective degradation-aware mechanism to distinguish different degradation remains challenging. AirNet [32] addresses this by employing contrastive learning to extract degradation representations, while HAIR [7] leverages a degradation-awareness classifier to capture global information and adaptively generate parameters via a hyper-selection network.
Recently, prompt learning [43, 31] has emerged as a promising paradigm for all-in-one image restoration due to its inherent flexibility and task-specific adaptability. PromptIR [52] integrates visual prompts for dynamic guidance, while PIP [36] refines them into degradation-aware and basic restoration prompts. To enable more controllable restoration, InstructIR [12] utilizes manually crafted instructions to guide the restoration process. Moreover, with the rise of VLMs, several studies [1, 44, 53] have explored the use of pre-trained VLMs to encode both textual and visual information for identifying degradation types. In the field of all-in-one HSI restoration, PromptHSI [30] introduces a text-to-feature modulation mechanism that applies textual prompts to composite degradation restoration in HSI, but it remains inherently limited to a single degradation mode. Overall, existing prompt-based all-in-one methods struggle to effectively characterize different degradation types in HSIs and lack guidance for spectral feature restoration.
Figure 3: (a) The architecture of the proposed MP-HSIR. (b) Prompt-Guided Spatial-Spectral Transformer Block (PGSSTB). (c) Design of the Prompt-Guided Spectral Self-Attention (PGSSA) Module. (d) Text-Visual Synergistic Prompt (TVSP) Module.
3 Method
In this section, we present the multi-prompt framework for universal HSI restoration. First, we review the task of HSI restoration in 3.1. Then, we describe the overall architecture of the MP-HSIR framework in 3.2. Finally, we introduce the Spectral Self-Attention with Prompt Guidance module and the Text-Visual Synergistic Prompts module in 3.3 and 3.4.
3.1 Preliminary: Hyperspectral Image Restoration
HSI restoration aims to recover the clean image $\mathcal{X}$ from the degraded observation $\mathcal{Y}$ [15, 77], modeled as:
$\mathcal{Y}=\mathcal{H}(\mathcal{X})+\mathcal{N},$ (1)
where $\mathcal{H}$ represents multiplicative degradation operations and $\mathcal{N}$ denotes additive degradation operators. For the denoising task, $\mathcal{H}$ is an identity operation. For the deblurring task, $\mathcal{\mathcal{H}}$ represents the blurring operation. For the inpainting task, $\mathcal{H}$ denotes a binary mask. In practice, the specific degradation in HSI often involves a combination of unknown multiplicative and additive operators. Without prior knowledge of the degradation type, selecting and deploying a specific restoration method requires maintaining multiple specialized models, inevitably leading to high storage costs and complex model management.
3.2 Overall Pipeline
Given the degraded observation $\mathcal{Y}\in\mathbb{R}^{H\times W\times B}$, where $H\times W$ represents the spatial resolution and $B$ denotes the number of spectral bands, MP-HSIR first applies a single-layer convolution to extract shallow features. As illustrated in Figure 11 (a), these features are processed through a three-level hierarchical encoder with downsampling operations. This process reduces spatial resolution while increasing channel capacity to obtain multi-scale deep features ${F}_{l}^{e}\in\mathbb{R}^{H\times W\times C}$, where $C$ is the number of channels. Each encoder-decoder pair consists of multiple Prompt-Guided Spatial-Spectral Transformer Blocks (PGSSTBs), with the number of blocks progressively increasing at deeper levels.
To integrate degradation information into the decoding process, we introduce the Text-Visual Synergistic Prompt (TVSP) module at each skip connection between the encoder and decoder. The TVSP module leverages textual and visual prompts to incorporate semantic guidance and fine-grained details, effectively capturing degradation characteristics for restoration. Finally, the restored HSI is generated via a $3\times 3$ convolution layer followed by an image-level residual connection. The following sections detail the designs of the PGSSTB and TVSP modules.
3.3 Prompt-Guided Spatial-Spectral Transformer
The PGSSTB is the core component of MP-HSIR, integrating Spatial Self-Attention (SSA), Prompt-Guided Spectral Self-Attention (PGSSA), and a gated multi-layer perceptron (GMLP) [38]. To enhance training stability and convergence speed, layer normalization and skip connections [20] are incorporated, as shown in Figure 11 (b). SSA computes attention along the spatial dimension, capturing the non-local similarity in HSI. To improve computational efficiency while preserving the ability to model long-range dependencies, a sliding window [40] of size $P\times P$ is adopted.
While SSA captures spatial dependencies, spectral self-attention models long-range correlations across spectral bands, which is crucial for exploiting HSI’s rich spectral information. To compute spectral similarity, the input features are transposed and reshaped into $F^{in}\in\mathbb{R}^{C\times HW}$ and processed through two complementary branches: global spectral self-attention and prompt-guided local spectral self-attention. The outputs from both branches are then fused and selectively refined using a GMLP. The following sections detail the two branches of PGSSA.
Global Spectral Self-Attention. Global spectral self-attention captures inter-channel dependencies across the entire HSI, as illustrated in the upper part of Figure 11 (c). Specifically, the input feature $F^{in}$ is first projected into the query $Q^{g}\in\mathbb{R}^{C\times HW}$, key $K^{g}\in\mathbb{R}^{C\times HW}$, and value $V^{g}\in\mathbb{R}^{C\times HW}$ using $1\times 1$ point-wise convolutions $W^{P}$, followed by $3\times 3$ depth-wise convolutions $W^{D}$:
$\left[{Q}^{g},{K}^{g},{V}^{g}\right]={W}^{D}{W}^{P}\left[{F}^{in},{F}^{in},{F}^{in}\right].$ (2)
To capture diverse spectral relationships across different bands, we further divide $Q^{g}$, $K^{g}$, and $V^{g}$ into multiple heads. Within each head, the global spectral attention map $A^{g}\in\mathbb{R}^{C\times C}$ is computed using the element-wise dot product between $Q^{g}$ and $K^{g}$, formulated as:
${A}^{g}=\text{Softmax}\left(\frac{{Q}^{g}\cdot{K}^{g}}{\epsilon}\right),$ (3)
$\text{Attention}\left({Q}^{g},{K}^{g},{V}^{g}\right)={W}^{P}\left({A}^{g}{V}^{g}\right),$
where $\epsilon$ is a learnable scaling factor that normalizes the results of the dot product.
Prompt-Guided Local Spectral Self-Attention. Unlike global spectral self-attention, local spectral self-attention models spectral channel dependencies within localized spatial regions, as illustrated in the lower part of Figure 11 (c). Specifically, the input feature $F^{in}$ is divided into $\frac{HW}{P^{2}}$ non-overlapping local patches $\{M_{1}^{in},M_{2}^{in},\dots,M_{J}^{in}\}$, where each patch $M_{j}^{in}\in\mathbb{R}^{C\times P^{2}}$ has the same spatial size as the spatial self-attention window. To extract representative spectral features $M_{j}^{s}\in\mathbb{R}^{1\times C}$ from each patch, global average pooling is applied along the spatial dimension:
$M_{j}^{s}=\left(\text{ AvgPool }\!\left(M_{j}^{in}\right)\right)^{\top}.$ (4)
However, due to the susceptibility of spectral structures to various degradations, purely local modeling may lead to an overly dispersed attention distribution, preventing the model from focusing on the most informative spectral channels. To address this, a learnable spectral prompt, $P_{S}\in\mathbb{R}^{L\times D}$, is introduced to guide the attention mechanism toward essential spectral patterns, where $L$ denotes the number of universal low-rank spectral patterns in $P_{S}$, and $D$ represents the dimension. Specifically, $M_{j}^{s}$ undergoes a linear transformation followed by Softmax normalization to produce weights. These weights are then used to compute a weighted sum of $P_{S}$, which is subsequently projected to obtain the query $Q^{l}$. Meanwhile, $M_{j}^{s}$ is projected into the same low-dimensional space and further transformed to generate the key $K^{l}$ and value $V^{l}$, formulated as:
$Q^{l}=\text{ Softmax }\!\left(M_{j}^{s}W_{1}^{l}\right)P_{S}W_{3}^{l},$ (5)
$\left[{K}^{l},{V}^{l}\right]=M_{j}^{s}W_{2}^{l}W_{3}^{l},$
where the linear layer weights $W_{1}^{l}$, $W_{2}^{l}$, and $W_{3}^{l}$ have dimensions $C\times L$, $C\times D$, and $D\times D$, respectively.
The spectral prompt $P_{S}$, learned during training, provides universally applicable low-rank spectral patterns for local spectral self-attention, enabling any local spectral feature to be represented as a linear combination of these patterns. This enhances the performance of local spectral modeling and facilitates spectral information recovery, thereby improving the model’s generalization across different datasets and degradation scenarios. Finally, the local spectral attention map ${A}^{l}\in\mathbb{R}^{D\times D}$ can be obtained by:
${A}^{l}=\text{Softmax}\left(\frac{{Q}^{l}\cdot{K}^{l}}{\epsilon}\right),$ (6)
$\text{Attention}\left({Q}^{l},{K}^{l},{V}^{l}\right)={W}^{P}\left({A}^{l}{V}^{l}\right),$
where $W^{P}$ remaps the low-dimensional spectral features to the original dimension $M_{j}^{out}\in\mathbb{R}^{C\times 1}$, which is then used to weight $M_{j}^{in}$ to obtain the final output.
3.4 Text-Visual Synergistic Prompt
The spectral prompt exhibits strong robustness to HSIs with varying degradations but lacks explicit guidance on the degradation type. To address this limitation, we introduce the TVSP module, as illustrated in Figure 11 (d), which integrates textual prompts and learnable visual prompts to provide degradation-specific information for the restoration network.
First, we employ a degradation predictor $\Phi$ [11] to classify the degradation type of the input observation (details are provided in Section A of the supplementary material). Based on predefined textual descriptions $T_{text}$ corresponding to different degradation types, a frozen CLIP model [55] generates embeddings that serve as the textual prompt $P_{T}$. To bridge the domain gap between RGB and HSI, these embeddings are transformed into learnable parameters, enabling the network to adapt and refine the textual prompts during training for improved alignment with HSI characteristics. The final textual prompt is selected based on the classification result:
$P_{T}=\Phi(\mathcal{X})\cdot\operatorname{Clip}\left(T_{text}\right).$ (7)
However, the textual prompt $P_{T}$ provides only global degradation information and lacks pixel-level precision. To address this, we introduce a learnable visual prompt $P_{V}$ and fuse it with the textual prompt via a cross-attention. This allows $P_{T}$ to regulate global degradation while $P_{V}$ refines local features, enabling degradation information to be effectively integrated and adjusted, formally expressed as:
$F_{l}^{out}=\operatorname{Concat}\left(F_{l}^{e},\text{Attention }\!\left(P_{T},P_{V}\right)\right),$ (8)
where the output $F_{l}^{out}$ is finally passed to the decoder via a skip connection. Compared to VLM-based dual prompts, TVSP is more flexible and adaptive, enabling prompt adjustments for different HSI domains.
4 Experiments
Type Method Gaussian Denoising (Sigma = 30, 50, 70) Method Complex Denoising (Case = 1, 2, 3, 4)
ICVL [3] ARAD [4] Xiong’an [72] ICVL [3] ARAD [4] WDC [81]
PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$
Task Specific QRNN3D [63] 39.99 / 0.947 39.18 / 0.932 36.06 / 0.829 QRNN3D [63] 41.60 / 0.966 41.11 / 0.960 30.82 / 0.868
SST [33] 41.31 / 0.959 40.86 / 0.955 37.53 / 0.850 SST [33] 42.43 / 0.971 41.95 / 0.967 32.71 / 0.889
SERT [34] 41.55 / 0.967 41.09 / 0.959 37.83 / 0.859 SERT [34] 43.31 / 0.976 42.87 / 0.973 33.31 / 0.903
LDERT [35] 41.92 / 0.969 41.47 / 0.965 38.14 / 0.865 LDERT [35] 43.42 / 0.977 43.02 / 0.974 33.49 / 0.904
All in One AirNet [32] 39.76 / 0.943 39.19 / 0.920 31.94 / 0.668 AirNet [32] 40.68 / 0.959 40.09 / 0.945 28.07 / 0.726
PromptIR [52] 40.25 / 0.953 39.69 / 0.945 32.99 / 0.684 PromptIR [52] 41.29 / 0.965 40.71 / 0.954 28.83 / 0.735
PIP [36] 40.87 / 0.958 40.25 / 0.952 32.64 / 0.674 PIP [36] 41.67 / 0.968 41.17 / 0.962 28.50 / 0.724
HAIR [7] 40.51 / 0.956 39.91 / 0.949 32.54 / 0.679 HAIR [7] 40.63 / 0.958 40.15 / 0.947 28.18 / 0.729
InstructIR [12] 41.02 / 0.960 40.32 / 0.954 31.74 / 0.666 InstructIR [12] 40.12 / 0.957 39.81 / 0.952 27.49 / 0.707
PromptHSI [30] 40.65 / 0.960 40.04 / 0.956 38.07 / 0.881 PromptHSI [30] 39.14 / 0.955 38.75 / 0.936 33.77 / 0.912
MP-HSIR (Ours) 41.62 / 0.964 41.09 / 0.960 38.81 / 0.897 MP-HSIR (Ours) 42.29 / 0.971 41.99 / 0.969 34.07 / 0.918
Type Method Gaussian Deblurring (Radius = 9, 15, 21 & 7, 11, 15) Method Super-Resolution (Scale = 2, 4, 8)
ICVL [3] PaviaC [24] Eagle [51] ARAD [4] PaviaU [24] Houston [66]
PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$
Task Specific Stripformer [58] 46.03 / 0.988 37.13 / 0.913 41.96 / 0.958 SNLSR [23] 36.05 / 0.898 30.55 / 0.730 31.70 / 0.787
FFTformer [27] 46.65 / 0.988 37.96 / 0.921 42.76 / 0.962 MAN [60] 36.88 / 0.911 30.92 / 0.733 32.03 / 0.791
LoFormer [47] 47.15 / 0.989 37.72 / 0.917 42.59 / 0.962 ESSAformer [78] 37.40 / 0.918 31.34 / 0.738 32.37 / 0.793
MLWNet [14] 47.66 / 0.990 39.01 / 0.928 44.04 / 0.969 SRFormer [80] 38.02 / 0.922 31.80 / 0.751 32.80 / 0.803
All in One AirNet [32] 47.21 / 0.989 37.76 / 0.918 42.42 / 0.963 AirNet [32] 36.73 / 0.910 30.93 / 0.735 32.00 / 0.791
PromptIR [52] 47.67 / 0.990 38.72 / 0.927 43.74 / 0.968 PromptIR [52] 37.37 / 0.918 31.53 / 0.746 32.73 / 0.799
PIP [36] 47.52 / 0.990 38.52 / 0.924 43.17 / 0.967 PIP [36] 38.36 / 0.926 31.77 / 0.749 32.89 / 0.805
HAIR [7] 46.45 / 0.988 37.93 / 0.919 42.77 / 0.964 HAIR [7] 36.84 / 0.915 31.47 / 0.745 32.54 / 0.796
InstructIR [12] 34.05 / 0.798 27.03 / 0.602 32.68 / 0.766 InstructIR [12] 36.51 / 0.913 31.34 / 0.741 32.52 / 0.793
PromptHSI [30] 32.52 / 0.817 36.73 / 0.912 39.75 / 0.954 PromptHSI [30] 35.03 / 0.904 30.75 / 0.732 31.85 / 0.775
MP-HSIR (Ours) 48.07 / 0.990 39.00 / 0.927 43.72 / 0.968 MP-HSIR (Ours) 38.25 / 0.924 31.97 / 0.760 33.06 / 0.810
Method Inpainting (Mask Rate = 0.7, 0.8, 0.9) Method Dehazing (Omega = 0.5, 0.75, 1) Method Completion (Rate = 0.1, 0.2 0.3)
ICVL [3] Chikusei [73] PaviaU [24] Eagle [51] ARAD [4] Berlin [49]
PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$
NAFNet [8] 44.39 / 0.987 39.82 / 0.953 SGNet [45] 34.28 / 0.963 37.22 / 0.976 NAFNet [8] 47.04 / 0.995 38.35 / 0.907
Restormer [76] 45.79 / 0.990 36.33 / 0.899 SCANet [18] 36.59 / 0.978 39.64 / 0.985 Restormer [76] 48.34 / 0.995 35.07 / 0.606
DDS2M [48] 42.18 / 0.969 34.94 / 0.887 MB-Taylor [54] 37.99 / 0.983 41.03 / 0.991 SwinIR [37] 49.75 / 0.995 35.45 / 0.886
HIR-Diff [50] 38.91 / 0.949 37.65 / 0.916 DCMPNet [79] 37.20 / 0.985 40.24 / 0.990 MambaIR [16] 50.26 / 0.995 36.12 / 0.888
AirNet [32] 42.60 / 0.981 37.46 / 0.919 AirNet [32] 35.59 / 0.965 38.83 / 0.981 AirNet [32] 45.27 / 0.992 35.91 / 0.624
PromptIR [52] 46.38 / 0.990 38.07 / 0.930 PromptIR [52] 37.41 / 0.982 40.73 / 0.992 PromptIR [52] 46.60 / 0.994 40.45 / 0.652
PIP [36] 43.37 / 0.982 38.43 / 0.930 PIP [36] 37.62 / 0.982 40.74 / 0.990 PIP [36] 47.36 / 0.993 37.20 / 0.668
HAIR [7] 44.02 / 0.982 38.05 / 0.927 HAIR [7] 36.76 / 0.978 40.86 / 0.992 HAIR [7] 45.08 / 0.992 38.01 / 0.650
InstructIR [12] 44.07 / 0.986 36.11 / 0.907 InstructIR [12] 34.72 / 0.971 37.65 / 0.982 InstructIR [12] 51.31 / 0.997 35.97 / 0.580
PromptHSI [30] 41.48 / 0.972 37.33 / 0.946 PromptHSI [30] 36.77 / 0.974 39.78 / 0.984 PromptHSI [30] 47.34 / 0.994 43.36 / 0.972
MP-HSIR (Ours) 51.53 / 0.996 43.63 / 0.979 MP-HSIR (Ours) 39.59 / 0.986 42.41 / 0.995 MP-HSIR (Ours) 56.48 / 0.999 49.50 / 0.987
Table 1: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods on 7 HSI restoration tasks. The best and second-best performances are highlighted in red and blue, respectively.
4.1 Experimental setup
Datasets. We employ 13 hyperspectral datasets for training and testing, comprising two natural scene HSI datasets and 11 remote sensing HSI datasets. The natural scene datasets include ICVL [3] and ARAD [4], while the remote sensing datasets consist of Xiong’an [72], WDC [81], Pavia City and Pavia University [24], Houston [66], Chikusei [73], Eagle [51], Berlin [49], APEX [25], Urban [5], and 10 EO-1 Hyperion scenes collected in this study. For the natural scene HSI datasets, 1,000 images are cropped into 64×64×31 patches for training, while an additional 100 non-overlapping images are reserved for testing. For the first eight remote sensing datasets, 80% of each image is cropped into 64×64×100 patches for training, with the remaining regions used for testing. The last three remote sensing datasets are designated for real-world experiments. Due to the substantial domain differences between natural scenes and remote sensing data, we train models separately on each dataset category. To ensure a comprehensive evaluation, multiple test datasets are randomly assigned to each task, ensuring that both natural scene and remote sensing HSI datasets are included in every task. Moreover, normalization is conducted on all datasets. A detailed description of the datasets is provided in Section B of the supplementary material.
Settings. To validate the effectiveness and generalization capability of the proposed model, we conducted evaluations under three different settings. (1) All-in-one: A unified model was trained on seven HSI restoration tasks, including Gaussian denoising, complex denoising, Gaussian deblurring, super-resolution, inpainting, dehazing, and band completion. (2) Generalization: Few-shot and zero-shot learning were performed on unseen tasks. Motion deblurring was evaluated after fine-tuning, while Poisson denoising was assessed directly using the pre-trained model. (3) Real-World: For real-world denoising, following the settings of [33], the model was fine-tuned on the APEX dataset [25] and evaluated on the Urban dataset [5]. For real-world dehazing, the model was directly evaluated on the EO-1 Hyperion data. Detailed experimental settings for each task, including degradation synthesis models and intensity ranges, are provided in Section C of the supplementary material.
Implementation Details. Due to the difference in data volume, we employed distinct training strategies for natural scene and remote sensing HSI datasets. Both utilized the AdamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$) with L1 loss and a batch size of 32. For natural HSIs, training was conducted for 100 epochs with an initial learning rate of $2\times 10^{-4}$, which was gradually reduced to $1\times 10^{-6}$ using cosine annealing [41]. In contrast, for remote sensing HSIs, the initial learning rate was set to $1\times 10^{-4}$, and training was extended to 300 epochs. Given the higher spectral dimensionality of remote sensing HSIs, the width of the all-in-one model was increased by a factor of 1.5, while task-specific models were fine-tuned to achieve optimal performance.
      Input / 14.07        SST [33] / 30.53      LDERT [35] / 30.74   PromptIR [52] / 26.38   InstructIR [12] / 24.63 PromptHSI [30] / 30.78      Ours / 31.36         GT / PSNR (dB)
      Input / 14.20       NAFNet [8] / 40.82    Restormer [76] / 42.16   PromptIR [52] / 42.65   InstructIR [12] / 41.43 PromptHSI [30] / 38.52      Ours / 47.67         GT / PSNR (dB)
      Input / 18.24     MB-Taylor [54] / 38.34 DCMPNet [79] / 37.68 PromptIR [52] / 37.35 InstructIR [12] / 33.76 PromptHSI [30] / 37.82      Ours / 39.49         GT / PSNR (dB)
Figure 4: [All-in-one] Visual comparison results on complex denoising, inpainting, and dehazing. The zoomed-in regions are highlighted with red boxes, the corresponding zoomed results are marked with green boxes, and the residual maps are outlined with yellow boxes.
Figure 5: The spectral errors of the restored HSIs across multiple tasks. (a) Gaussian denoising. (b) Complex denoising. (c) Gaussian deblurring. (d) Super-resolution. (e) Inpainting. (f) Dehazing. (g) Band completion. (h) Poisson denoising. Zoom in for the best view.
Type Method Motion Deblurring Methods Poisson Denoising
 PSNR $\uparrow$  SSIM $\uparrow$  PSNR $\uparrow$  SSIM $\uparrow$
Task Specific Stripformer [58]  37.91  0.948 QRNN3D [63]  35.98  0.910
FFTformer [27]  38.42  0.950 SST [33]  37.51  0.914
LoFormer [47]  38.94  0.953 SERT [34]  37.46  0.913
MLWNet [14] 39.28 0.959 LDERT [35] 37.92 0.915
All in One AirNet [32]  37.43  0.947 AirNet [32]  33.91  0.896
PromptIR [52]  38.77  0.951 PromptIR [52]  34.83  0.902
PIP [36]  38.55  0.950 PIP [36]  34.52  0.901
HAIR [7]  37.58  0.948 HAIR [7]  34.38  0.901
InstructIR [12]  36.58  0.942 InstructIR [12]  33.96  0.895
PromptHSI [30]  36.79  0.945 PromptHSI [30]  37.25  0.912
MP-HSIR (Ours) 40.06 0.965 MP-HSIR (Ours) 38.56 0.922
Table 2: [Generalization] The results of motion deblurring on the ICVL dataset [3] and Poisson denoising on the Houston dataset [66]. Task-specific methods are trained on the entire dataset, while all-in-one methods are fine-tuned using only 5% of the data.
Evaluation Metrics and Comparisons Methods. We employed PSNR and SSIM [61] as quantitative metrics for evaluation. The comparison methods included six all-in-one models, i.e., AirNet [32], PromptIR [52], PIP [36], HAIR [7], InstructIR [12], and PromptHSI [30], along with 22 state-of-the-art task-specific models.
4.2 Comparison with state-of-the-art methods
All-in-one. Table 1 presents the quantitative performance of seven all-in-one methods and 22 task-specific approaches across seven HSI restoration tasks. Our method consistently outperforms all all-in-one methods and surpasses state-of-the-art task-specific approaches in multiple tasks, achieving significant PSNR improvements in inpainting, dehazing, and band completion. Furthermore, our method attains the best or near-best performance across all hyperspectral remote sensing datasets. Notably, InstructIR and PromptHSI exhibit anomalous results in the deblurring task due to their inability to effectively distinguish different types of HSI degradation. Visualization comparison of prompts is provided in Section D of the supplementary material.
Figure 9 presents visual comparisons for complex denoising, inpainting, and dehazing. The residual maps highlight the superiority of our method in structural recovery and fine-detail reconstruction. Figure 12 illustrates the normalized digital number (DN) error between the restored and ground truth images across spectral bands. Leveraging spectral prompts, our method achieves the most accurate spectral reconstruction across all tasks.
        Input               SST [33]            LDERT [35]           AirNet [32]          PromptIR [52]        InstructIR [12]        PromptHSI [30]           Ours
        Input            MB-Taylor [54]        DCMPNet [79]          AirNet [32]          PromptIR [52]        InstructIR [12]        PromptHSI [30]           Ours
Figure 6: [Real-world] Visual comparison results on real denoising and real dehazing. Zoom in for the best view.
Method PSNR $\uparrow$ SSIM $\uparrow$ Params (M)
Baseline (Only Spatial SA) 39.24 0.963 20.93
+ Textual Prompt $P_{T}$ 39.62 0.964 21.51
+ Visual Prompt $P_{V}$ 39.57 0.964 23.68
+ Textual Prompt $P_{T}$ + Visual Prompt $P_{V}$ 39.90 0.964 24.26
+ Global Spectral SA + $P_{T}$ + $P_{T}$ 40.63 0.969 30.07
+ Local Spectral SA + $P_{T}$ + $P_{V}$ 40.46 0.968 24.43
+ Local Spectral SA + $P_{T}$ + $P_{V}$ + Spectral Prompt $P_{S}$ 41.05 0.971 25.10
Full Model 41.98 0.974 30.91
Table 3: Ablation study to verify the effectiveness of modules on Chikusei dataset in inpainting task with mask ratio = 0.9.
Generalization. Table 2 presents the experimental results for motion blur removal using few-shot learning and Poisson noise removal using zero-shot learning. The proposed method exhibits superior generalization ability compared to existing approaches under both learning paradigms.
Real-world. As shown in Figure 13, we present a comparative visualization of all-in-one and task-specific methods applied to real-world denoising and dehazing scenarios. The results highlight the proposed method’s effectiveness in restoring texture details and spectral structure, while mitigating common artifacts like over-smoothing and spectral distortion.
Additional Results. More quantitative results and visualizations are in Section E of the supplementary material.
4.3 Model Analysis
Ablation Study. To assess the effectiveness of each module in the proposed MP-HSIR framework, we conducted an ablation study, starting with a baseline model that includes only spatial self-attention (SA) and then progressively incorporating textual prompts $P_{T}$, learnable visual prompts $P_{V}$, global spectral self-attention, local spectral self-attention, and spectral prompts $P_{S}$. As shown in Table 3, both accuracy metrics for the inpainting task on the Chikusei dataset progressively improve with the addition of each module. Notably, the local spectral self-attention module guided by spectral prompts provides the most significant performance gain, attributed to the additional spectral guidance offered by spectral prompts. Results for other tasks are provided in Section E of the supplementary material.
Visualization of Spectral Prompt. To further investigate the role of spectral prompts, we visualize the activation distribution of universal low-rank spectral patterns under both noise and blur degradations in Figure 14. By selecting identical regions from both degraded images for comparison, we observe that despite different degradation types, the activation patterns remain highly consistent within the same regions. This indicates that spectral prompts effectively capture key spectral information in HSI and demonstrate strong robustness against different types of degradation.
Figure 7: Visualization of the activation proportions for universal low-rank spectral patterns within the spectral prompt.
5 Conclusion
HSI restoration is challenging due to diverse and often unknown degradation types, while existing methods are typically limited to specific degradation scenarios. In this work, we propose MP-HSIR, a universal HSI restoration framework that integrates spectral, textual, and visual prompts. MP-HSIR introduces a novel prompt-guided spatial-spectral transformer, where spectral prompts provide universal low-rank spectral patterns to enhance spectral reconstruction. Furthermore, the text-visual synergistic prompts combine high-level semantic representations with fine-grained visual features to guide degradation modeling. Extensive experiments on 9 HSI restoration tasks—including all-in-one restoration, generalization testing, and real-world scenarios—demonstrate the superior robustness and adaptability of MP-HSIR under diverse degradation conditions.
Supplementary Materials
F Degradation Preictor
The proposed universal multi-prompt framework, MP-HSIR, can perform specific restoration tasks either based on human instructions or autonomously via the degradation predictor. In this work, we employ a ResNet-34 network with fast Fourier convolution [11] to train classification models separately on natural scene and remote sensing hyperspectral datasets. The binary cross-entropy loss used for training is defined as follows:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log p_{i}+\left(1-y_{i}\right)\log\left(1-p_{i}\right)\right],$ (9)
where $N$ denotes the number of samples, $p_{i}$ represents the output after sigmoid activation, and $y_{i}$ is the corresponding ground-truth label. The training was conducted with a batch size of 64, using the same optimizer as in the restoration experiments. The initial learning rate was set to $1\times 10^{-4}$ and progressively decreased to $1\times 10^{-6}$ via cosine annealing [41]. The model was trained for 1000 epochs on the natural scene hyperspectral dataset and 4000 epochs on the remote sensing hyperspectral dataset.
Table 4 presents the accuracy and precision of the degradation predictor for both hyperspectral datasets. The results show that the predictor achieves 100% accuracy and precision across all tasks, demonstrating MPIR-HSI’s effectiveness in supporting all trained blind restoration tasks.
G Dataset Details
This section provides a comprehensive overview of the 13 datasets used across 9 hyperspectral image (HSI) restoration tasks and real-world scenarios, as summarized in Table 5.
ARAD [4]. The ARAD dataset, derived from the NTIRE 2022 Spectral Recovery Challenge, was collected using a Specim IQ hyperspectral camera. It consists of 1,000 images, with 900 allocated for training and 50 for testing.
ICVL [3]. The ICVL dataset was obtained using a Specim PS Kappa DX4 hyperspectral camera combined with a rotating stage for spatial scanning. It contains 201 images, with 100 for training and 50 for testing, ensuring no scene overlap.
Xiong’an [72]. The Xiong’an dataset was captured using an imaging spectrometer developed by the Chinese Academy of Sciences. Three central regions of size 512 × 512 were randomly cropped for testing, while the remaining areas were used for training.
Task Natural Scene Remote Sensing
Accuracy $\uparrow$ Precision $\uparrow$ Accuracy $\uparrow$ Precision $\uparrow$
Gaussian Denoising 100.00 100.00 100.00 100.00
Complex Denoising 100.00 100.00 100.00 100.00
Gaussian Deblurring 100.00 100.00 100.00 100.00
Super-Resolution 100.00 100.00 100.00 100.00
Inpainting 100.00 100.00 100.00 100.00
Dehazing 100.00 100.00 100.00 100.00
Band Completion 100.00 100.00 100.00 100.00
Table 4: Accuracy and precision results of the degradation predictor in degradation classification
[…]
Table 5: Properties of 13 Natural Scene and Remote Sensing Hyperspectral Datasets.
WDC [81]. The Washington DC (WDC) dataset was captured by a Hydice sensor. A central region of size 256 × 256 was selected for testing, with the remainder used for training.
PaviaC [24]. The Pavia Center (PaviaC) dataset was acquired using a ROSIS sensor, following the same partitioning strategy as the WDC dataset.
PaviaU [24]. The Pavia University (PaviaU) dataset was also collected using a ROSIS sensor, with the same partitioning strategy as WDC.
Houston [66]. The Houston dataset was obtained using an ITRES CASI-1500 sensor, employing the same partitioning strategy as WDC.
Chikusei [73]. The Chikusei dataset was captured using an HH-VNIR-C sensor. Four 512 × 512 regions were randomly cropped for testing, with the remaining areas used for training.
Eagle [51]. The Eagle dataset was collected using an AsiaEAGLE II sensor, following the same partitioning strategy as WDC.
Berlin [49]. The Berlin dataset utilizes only the HyMap image from the BerlinUrbGrad dataset. A 512 × 512 central region was randomly cropped for testing, while the remaining data were used for training.
Urban [5]. The Urban dataset, collected using a Hydice sensor, is specifically used for real-world denoising experiments.
APEX [25]. The APEX dataset exhibits characteristics similar to the Urban dataset and is primarily used for fine-tuning pre-trained models.
EO-1 [13]. The EO-1 dataset was captured by the Hyperion sensor. Ten scenes were collected for testing, with 67 invalid bands removed, retaining 175 valid bands for real-world dehazing experiments.
All datasets underwent min-max normalization, and training samples were uniformly cropped to 64 × 64.
H Detailed Experimental Setup
In this section, we provide a detailed description of the experimental settings for the 9 HSI restoration tasks.
Gaussian Denoising. Each image was corrupted by zero-mean independent and identically distributed (i.i.d.) Gaussian noise with sigma ranging from 30 to 70. For testing, sigma = 30, 50, and 70 were selected for evaluation.
Complex Denoising. Each image was corrupted with one of the following four noise scenarios:
1) Case 1 (Non-i.i.d. Gaussian Noise): All bands were corrupted by non-i.i.d. Gaussian noise with standard deviations randomly selected from 10 to 70.
2) Case 2 (Gaussian Noise + Stripe Noise): All bands were corrupted by non-i.i.d. Gaussian noise, and one-third of the bands were randomly selected to add column stripe noise with intensities ranging from 5% to 15%.
3) Case 3 (Gaussian Noise + Deadline Noise): The noise generation process was similar to Case 2, but stripe noise was replaced by deadline noise.
4) Case 4 (Gaussian Noise + Impulse Noise): All bands were corrupted by non-i.i.d. Gaussian noise, and one-third of the bands were randomly selected to add impulse noise with intensities ranging from 10% to 70%.
Gaussian Deblurring. An empirical formula was used to calculate the standard deviation $\sigma$ based on the Gaussian kernel size $K_{S}$, formulated as:
$\sigma=0.3\times\left(\frac{K_{S}-1}{2}-1\right)+0.8.$ (10)
For natural hyperspectral datasets, $K_{S}$ was set to 9, 15, and 21, while for remote sensing hyperspectral datasets, $K_{S}$ was set to 7, 11, and 15.
Super-Resolution. Bicubic interpolation was used to downsample the images, with downscaling factors of 2, 4, and 8. To ensure that the input and output image sizes of the all-in-one model remained consistent, an unpooling operation was applied to resize the downsampled HSIs to their original dimensions.
Inpainting. Random masks with rates of 0.7, 0.8, and 0.9 were applied to each image for the inpainting task.
Dehazing. To realistically simulate haze contamination, the haze synthesis method from [17] was adopted. Specifically, 100 haze masks were extracted from the cirrus band of Landsat-8 OLI and superimposed onto the original image according to the wavelength ratio to generate haze-affected HSIs, modeled as:
$I_{i}=J_{i}e^{\left(\frac{\lambda_{1}}{\lambda_{i}}\right)^{\gamma}\ln{t}_{1}}+A\left(1-e^{\left(\frac{\lambda_{1}}{\lambda_{i}}\right)^{\gamma}\ln{t}_{1}}\right),$ (11)
where $I$ is the hazy HSI, $J$ is the clear HSI, $A$ is the global atmospheric light, $\lambda$ is the wavelength, and $\gamma$ is the spatial function, which is set to 1. The reference transmission map ${t}_{1}$ is calculated from the cirrus band reflectance:
$t_{1}=1-\omega B_{9},$ (12)
where $\omega$ is a weighting factor controlling the haze intensity, and $B_{9}$ is the cirrus band reflectance. In the experiments, $\omega$ was set to [0.5, 0.75, 1], corresponding to different levels of haze contamination.
Band Completion. A certain proportion of bands were discarded for each image, with discard rates of 0.1, 0.2, and 0.3. The experimental results were evaluated only on the missing bands.
Motion Deblurring. The pre-trained model was fine-tuned and tested on this task with a blur kernel radius of 15 and a blur angle of 45 degrees.
Poisson Denoising. The pre-trained model was directly tested on this task with a Poisson noise intensity scaling factor of 10.
[…]
SST [33] 43.32 / 0.976 41.09 / 0.952 39.51 / 0.949 43.02 / 0.972 40.58 / 0.951 38.99 / 0.941 39.26 / 0.878 37.34 / 0.848 35.99 / 0.824
SERT [34] 43.53 / 0.978 41.32 / 0.966 39.82 / 0.956 43.21 / 0.975 40.84 / 0.959 39.21 / 0.945 39.54 / 0.885 37.58 / 0.859 36.37 / 0.833
[…]
PromptIR [52] 42.40 / 0.971 40.14 / 0.954 38.20 / 0.934 41.84 / 0.967 39.55 / 0.947 37.67 / 0.921 34.90 / 0.715 32.76 / 0.680 31.31 / 0.657
PIP [36] 43.00 / 0.974 40.69 / 0.958 38.94 / 0.941 42.33 / 0.970 40.07 / 0.953 38.36 / 0.933 34.51 / 0.704 32.43 / 0.671 30.98 / 0.647
HAIR [7] 42.53 / 0.972 40.23 / 0.957 38.78 / 0.939 42.03 / 0.968 39.76 / 0.950 37.95 / 0.928 34.51 / 0.712 32.22 / 0.675 30.89 / 0.650
InstructIR [12] 42.99 / 0.974 40.84 / 0.960 39.23 / 0.946 42.21 / 0.970 40.16 / 0.955 38.60 / 0.938 33.79 / 0.703 31.47 / 0.662 29.96 / 0.633
PromptHSI [30] 42.61 / 0.976 40.27 / 0.960 39.08 / 0.945 41.90 / 0.971 39.84 / 0.959 38.37 / 0.938 39.54 / 0.902 37.80 / 0.877 36.87 / 0.864
[…]
Table 6: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different Gaussian noise levels on Gaussian denoising tasks. The best and second-best performances are highlighted in red and blue, respectively.
[…]
SST [33] 43.38 / 0.976 42.69 / 0.973 42.51 / 0.972 41.16 / 0.964 42.84 / 0.973 42.38 / 0.971 42.01 / 0.970 40.56 / 0.956 33.85 / 0.907 33.69 / 0.905 33.37 / 0.901 29.92 / 0.841
SERT [34] 43.96 / 0.980 43.48 / 0.978 43.45 / 0.977 42.37 / 0.969 43.56 / 0.978 43.19 / 0.976 42.88 / 0.974 41.85 / 0.963 34.48 / 0.922 34.26 / 0.920 33.98 / 0.915 30.53 / 0.854
[…]
PromptIR [52] 42.76 / 0.973 41.93 / 0.969 41.43 / 0.969 39.04 / 0.947 42.26 / 0.969 41.54 / 0.964 40.90 / 0.964 38.12 / 0.919 29.84 / 0.761 29.71 / 0.754 29.39 / 0.747 26.38 / 0.676
PIP [36] 42.96 / 0.974 42.13 / 0.970 41.38 / 0.969 40.19 / 0.959 42.39 / 0.971 41.71 / 0.966 41.11 / 0.966 39.45 / 0.943 29.57 / 0.751 29.31 / 0.745 29.17 / 0.740 25.95 / 0.658
HAIR [7] 41.78 / 0.965 41.47 / 0.965 40.68 / 0.958 38.58 / 0.943 41.19 / 0.959 40.95 / 0.961 40.53 / 0.959 37.92 / 0.909 29.38 / 0.756 29.22 / 0.750 28.73 / 0.744 25.40 / 0.664
InstructIR [12] 41.29 / 0.963 40.89 / 0.961 39.94 / 0.958 38.46 / 0.945 40.72 / 0.960 40.38 / 0.957 39.94 / 0.956 38.21 / 0.934 28.66 / 0.736 28.47 / 0.730 28.19 / 0.725 24.63 / 0.637
PromptHSI [30] 40.61 / 0.967 40.36 / 0.965 39.30 / 0.960 36.27 / 0.927 40.22 / 0.951 39.98 / 0.955 39.42 / 0.953 35.39 / 0.884 34.93 / 0.931 34.87 / 0.930 34.49 / 0.928 30.78 / 0.857
[…]
Table 7: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different cases on Complex denoising tasks. The best and second-best performances are highlighted in red and blue, respectively.
Figure 8: Similarity matrices of prompt vector for seven HSI restoration tasks, with task names as in manuscript.
I Visualization Results of Prompts
In Table 8, we present the similarity matrices of prompt vectors for different degradation types, comparing PromptIR [52], InstructIR [12], and the proposed method. PromptIR relies exclusively on visual prompts, InstructIR utilizes textual prompts for guidance, and the proposed method integrates text-visual synergistic prompts to enhance degradation modeling. As illustrated in the figure, both PromptIR and InstructIR struggle to effectively differentiate between various HSI degradation types, whereas our method demonstrates distinct separation for each individual task, highlighting its superior capability in handling diverse degradation scenarios.
[…]
FFTformer [27] 49.83 / 0.994 46.43 / 0.989 43.69 / 0.981 39.82 / 0.942 37.63 / 0.919 36.44 / 0.901 45.51 / 0.980 42.18 / 0.962 40.59 / 0.944
LoFormer [47] 50.35 / 0.995 46.94 / 0.991 44.15 / 0.983 39.54 / 0.941 37.41 / 0.918 36.21 / 0.892 45.32 / 0.979 42.03 / 0.961 40.42 / 0.946
[…]
PromptIR [52] 51.00 / 0.996 47.70 / 0.992 44.32 / 0.983 40.41 / 0.948 38.62 / 0.929 37.13 / 0.904 46.17 / 0.981 43.45 / 0.968 41.60 / 0.955
PIP [36] 51.05 / 0.996 47.29 / 0.991 44.21 / 0.983 40.31 / 0.944 38.51 / 0.926 36.73 / 0.902 45.90 / 0.982 42.75 / 0.968 40.86 / 0.951
HAIR [7] 49.43 / 0.994 46.06 / 0.989 43.87 / 0.981 39.97 / 0.943 38.18 / 0.923 35.64 / 0.891 45.14 / 0.979 42.69 / 0.966 40.48 / 0.947
InstructIR [12] 24.19 / 0.533 33.44 / 0.878 44.51 / 0.983 19.66 / 0.329 25.12 / 0.583 36.32 / 0.894 23.35 / 0.477 34.05 / 0.872 40.64 / 0.949
PromptHSI [30] 25.16 / 0.619 30.89 / 0.852 41.52 / 0.980 38.73 / 0.936 36.63 / 0.913 34.84 / 0.888 42.45 / 0.972 39.34 / 0.952 37.46 / 0.938
[…]
Table 8: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different blur kernel radius on Gaussian deblurring tasks. The best and second-best performances are highlighted in red and blue, respectively.
J More Experimental Results
In this section, we present additional experimental results, including the results of more quantitative results, model efficiency, more results of ablation study, controllability analysis, and more visual comparisons.
[…]
MAN [60] 44.81 / 0.985 35.35 / 0.912 30.49 / 0.830 34.92 / 0.872 30.26 / 0.723 27.59 / 0.604 35.26 / 0.911 31.68 / 0.785 29.15 / 0.677
ESSAformer [78] 45.32 / 0.988 36.02 / 0.927 30.85 / 0.838 35.47 / 0.879 30.60 / 0.728 27.96 / 0.606 35.60 / 0.913 31.94 / 0.787 29.57 / 0.679
[…]
PromptIR [52] 45.33 / 0.988 36.00 / 0.927 30.77 / 0.838 35.57 / 0.883 30.81 / 0.735 28.20 / 0.619 35.96 / 0.915 32.38 / 0.798 29.85 / 0.684
PIP [36] 46.01 / 0.989 37.34 / 0.939 31.73 / 0.853 35.71 / 0.885 31.02 / 0.744 28.20 / 0.618 36.10 / 0.917 32.55 / 0.806 30.02 / 0.692
HAIR [7] 43.77 / 0.984 35.89 / 0.924 30.87 / 0.836 35.49 / 0.882 30.79 / 0.736 28.13 / 0.616 35.81 / 0.913 32.17 / 0.795 29.64 / 0.680
InstructIR [12] 43.47 / 0.984 35.46 / 0.921 30.61 / 0.834 35.25 / 0.879 30.71 / 0.732 28.05 / 0.613 35.68 / 0.909 32.15 / 0.789 29.73 / 0.681
PromptHSI [30] 40.25 / 0.975 35.41 / 0.931 29.35 / 0.806 34.84 / 0.871 30.13 / 0.722 27.27 / 0.602 35.34 / 0.912 31.62 / 0.778 28.59 / 0.635
[…]
Table 9: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different downsampling scales on Super-Resolution tasks. The best and second-best performances are highlighted in red and blue, respectively.
Dataset PromptIR InstructIR PromptHSI MP-HSIR
Urban 14.95 15.56 12.34 11.42
EO-1 17.99 19.71 18.13 16.54
Table 10: No-reference quality assessment on real datasets.
[…]
Restormer [76] 46.51 / 0.991 46.00 / 0.991 44.85 / 0.988 36.52 / 0.902 36.34 / 0.892 36.13 / 0.903
DDS2M [48] 45.41 / 0.989 43.34 / 0.983 37.80 / 0.935 36.77 / 0.906 35.23 / 0.901 32.83 / 0.854
[…]
PromptIR [52] 46.96 / 0.992 46.93 / 0.992 45.24 / 0.988 38.86 / 0.925 38.30 / 0.931 37.05 / 0.934
PIP [36] 44.37 / 0.985 43.47 / 0.983 42.26 / 0.978 38.74 / 0.922 38.58 / 0.930 37.98 / 0.938
HAIR [7] 44.83 / 0.983 44.30 / 0.983 42.92 / 0.981 38.43 / 0.921 38.28 / 0.928 37.43 / 0.932
InstructIR [12] 44.85 / 0.989 44.29 / 0.987 43.08 / 0.983 36.30 / 0.904 36.18 / 0.908 35.84 / 0.909
PromptHSI [30] 42.83 / 0.983 41.72 / 0.976 39.89 / 0.956 38.99 / 0.966 37.64 / 0.952 35.35 / 0.920
[…]
Table 11: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different mask rates on Inpainting tasks. The best and second-best performances are highlighted in red and blue, respectively.
[…]
SCANet [18] 39.01 / 0.986 36.54 / 0.978 34.21 / 0.969 41.92 / 0.991 39.68 / 0.987 37.31 / 0.978
MB-Taylor [54] 40.51 / 0.991 38.03 / 0.984 35.44 / 0.975 43.76 / 0.995 40.97 / 0.992 38.36 / 0.986
[…]
PromptIR [52] 40.34 / 0.991 37.43 / 0.983 34.47 / 0.971 43.55 / 0.995 40.69 / 0.992 37.94 / 0.988
PIP [36] 40.30 / 0.991 37.64 / 0.983 34.93 / 0.971 43.21 / 0.994 40.95 / 0.991 38.07 / 0.985
HAIR [7] 39.47 / 0.989 36.79 / 0.980 34.02 / 0.965 43.38 / 0.995 40.54 / 0.992 38.67 / 0.988
InstructIR [12] 38.24 / 0.986 34.57 / 0.974 31.36 / 0.954 40.90 / 0.991 38.07 / 0.985 33.99 / 0.971
PromptHSI [30] 38.62 / 0.982 36.48 / 0.975 35.22 / 0.964 40.88 / 0.986 40.49 / 0.984 37.98 / 0.981
[…]
Table 12: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different haze levels on Dehazing tasks. The best and second-best performances are highlighted in red and blue, respectively.
[…]
Restormer [76] 49.24 / 0.997 48.29 / 0.995 47.49 / 0.993 34.98 / 0.605 35.24 / 0.606 35.00 / 0.607
SwinIR [37] 50.95 / 0.997 49.80 / 0.995 48.49 / 0.993 36.88 / 0.950 34.69 / 0.855 34.78 / 0.854
[…]
PromptIR [52] 47.68 / 0.996 46.71 / 0.994 45.41 / 0.992 42.81 / 0.707 39.55 / 0.910 39.00 / 0.640
PIP [36] 48.35 / 0.995 47.37 / 0.994 46.37 / 0.991 38.60 / 0.706 36.58 / 0.657 36.43 / 0.641
HAIR [7] 46.27 / 0.994 44.92 / 0.992 44.04 / 0.990 40.04 / 0.705 36.54 / 0.607 37.45 / 0.639
InstructIR [12] 52.66 / 0.998 51.37 / 0.997 49.90 / 0.996 36.17 / 0.606 35.33 / 0.559 36.40 / 0.576
PromptHSI [30] 49.05 / 0.996 47.09 / 0.993 45.89 / 0.992 47.11 / 0.997 43.14 / 0.973 39.82 / 0.956
[…]
Table 13: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different mask rates on Band Completion tasks. The best and second-best performances are highlighted in red and blue, respectively.
J.1 More Quantitative Results
In this section, we provide a detailed quantitative comparison for 7 all-in-one HSI restoration tasks. The experimental results across different degradation levels are systematically presented for each task, including Gaussian denoising in Table 6, complex denoising in Table 7, Gaussian deblurring in Table 8, super-resolution in Table 9, inpainting in Table 11, dehazing in Table 12, and band completion in Table 13.
Figure 9: Progressive results (ours) and final results (others).
In addition, we further evaluate the performance on real-world hyperspectral data using the no-reference metric QSFL [71]. The results on the Urban and EO-1 datasets are reported in Table 10, providing a more comprehensive assessment of generalization ability under practical conditions.
[…]
PromptIR [52] 33.00 10.03 72.60 22.21
PIP [36] 27.80 10.66 58.26 22.08
HAIR [7] 7.68 2.72 17.28 6.46
InstructIR [12] 68.82 2.81 154.03 6.57
PromptHSI [30] 25.90 10.10 50.89 21.91
[…]
Table 14: Model complexity comparisons
Metric PromptIR InstructIR PromptHSI MP-HSIR
Inference Time (s) 0.087 0.065 0.282 0.083
Table 15: Inference time for remote sensing scenes (64×64×100).
Method PSNR $\uparrow$ SSIM $\uparrow$ Params (M)
Baseline (Only Spatial SA) 33.78 0.782 20.93
+ Textual Prompt $P_{T}$ 34.53 0.807 21.51
+ Visual Prompt $P_{V}$ 34.47 0.805 23.68
+ Textual Prompt $P_{T}$ + Visual Prompt $P_{V}$ 34.92 0.822 24.26
+ Global Spectral SA + $P_{T}$ + $P_{T}$ 35.20 0.835 30.07
+ Local Spectral SA + $P_{T}$ + $P_{V}$ 35.82 0.846 24.43
+ Local Spectral SA + $P_{T}$ + $P_{V}$ + Spectral Prompt $P_{S}$ 36.67 0.863 25.10
Full Model 37.17 0.874 30.91
Table 16: Ablation study to verify the effectiveness of modules on Xiong’an dataset in Gaussian denoising task with sigma = 70.
J.2 Model Efficiency
In this section, we present the parameter counts and computational costs of the all-in-one models for both natural scene and remote sensing hyperspectral datasets. Notably, the network width for remote sensing datasets is 1.5 times greater than that for natural scene datasets across all models. As demonstrated in Table 14, our method achieves a lower parameter count while maintaining competitive computational efficiency.
Furthermore, to evaluate the practical inference performance, Table 15 reports the average inference time of each method on remote sensing datasets. Our method maintains a favorable trade-off between efficiency and accuracy, demonstrating its suitability for large-scale deployment.
J.3 More results of Ablation Study
In this section, we present ablation studies on the textual prompts $P_{T}$, learnable visual prompts $P_{V}$, global spectral self-attention, local spectral self-attention, and spectral prompts $P_{S}$ across multiple tasks, as shown in Tables 16, 17, and 18. Overall, the addition of each module progressively improves the two accuracy metrics across all tasks.
Method PSNR $\uparrow$ SSIM $\uparrow$ Params (M)
Baseline (Only Spatial SA) 30.10 0.788 20.93
+ Textual Prompt $P_{T}$ 37.64 0.928 21.51
+ Visual Prompt $P_{V}$ 37.42 0.925 23.68
+ Textual Prompt $P_{T}$ + Visual Prompt $P_{V}$ 38.91 0.942 24.26
+ Global Spectral SA + $P_{T}$ + $P_{T}$ 39.20 0.944 30.07
+ Local Spectral SA + $P_{T}$ + $P_{V}$ 39.79 0.948 24.43
+ Local Spectral SA + $P_{T}$ + $P_{V}$ + Spectral Prompt $P_{S}$ 40.51 0.950 25.10
Full Model 41.36 0.952 30.91
Table 17: Ablation study to verify the effectiveness of modules on Eagle dataset in Gaussian deblurring task with radius = 15.
Method PSNR $\uparrow$ SSIM $\uparrow$ Params (M)
Baseline (Only Spatial SA) 32.52 0.967 20.93
+ Textual Prompt $P_{T}$ 34.24 0.965 21.51
+ Visual Prompt $P_{V}$ 34.13 0.964 23.68
+ Textual Prompt $P_{T}$ + Visual Prompt $P_{V}$ 34.92 0.969 24.26
+ Global Spectral SA + $P_{T}$ + $P_{T}$ 35.53 0.973 30.07
+ Local Spectral SA + $P_{T}$ + $P_{V}$ 35.64 0.974 24.43
+ Local Spectral SA + $P_{T}$ + $P_{V}$ + Spectral Prompt $P_{S}$ 36.13 0.976 25.10
Full Model 36.68 0.978 30.91
Table 18: Ablation study to verify the effectiveness of modules on PaviaU dataset in Dehazing task with Omega = 1.0.
Figure 10: Controllable Reconstruction: Removing Gaussian from Gaussian Noise + Gaussian Blur Degradation.
Figure 11: Controllable Reconstruction: Removing Complex Noise from Complex Noise + Gaussian Blur Degradation.
J.4 Controllable Results
In this section, we demonstrate the controllable restoration capability of the proposed method. Specifically, we conduct controlled restoration tasks under two composite degradation scenarios: Gaussian noise with Gaussian blur and complex noise with Gaussian blur, aiming to remove noise while preserving blur. As illustrated in Figures 10 and 11, our method can precisely remove specific degradation types through accurate guidance from textual prompts, highlighting its controllability and interpretability.
Building on this controllable restoration paradigm, we further explore sequential degradation removal, where multiple degradations are addressed step by step under prompt guidance. As shown in Figure 9, our method achieves superior results in this more challenging setting, outperforming other approaches in both visual quality and flexibility.
Figure 12: Visual comparison results of Gaussian denoising, complex denoising, Gaussian deblurring, and super-resolution, including the corresponding residual maps. Task-Specific represents the optimal task-specific method.
Figure 13: Visual comparison results of Inpainting, Dehazing, Band Completion, and Motion Deblurring, including the corresponding residual maps. Task-Specific represents the optimal task-specific method.
Figure 14: Visual comparison results of Poisson Denoising, including the corresponding residual maps. Task-Specific represents the optimal task-specific method.
Figure 15: Visual comparison results of Real Dehazing. Task-Specific represents the optimal task-specific method.
J.5 More Visual Results
In this section, we present further visual results for each task, including all-in-one experiments, generalization testing, and real-world scenarios. As shown in Figures 12, 13, 14, and 15, the visualization results indicate that our method achieves the best performance in restoring texture details and structural features.

### v3

@@
 Table 4: Accuracy and precision results of the degradation predictor in degradation classification
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Type Dataset Sensor Wavelength (nm) Channels Size GSD (m)
+Natural HSIs ARAD_1K Specim IQ 400–700 31 482×512 /
+ICVL Specim PS Kappa DX4 400–700 31 1392×1300 /
+Remote Sensing HSIs Xiong’an Unknown 400–1000 256 3750×1580 0.5
+WDC Hydice 400–2400 191 1208×307 5
+PaviaC ROSIS 430–860 102 1096×715 1.3
+PaviaU ROSIS 430–860 103 610×340 1.3
+Houston ITRES CASI-1500 364–1046 144 349×1905 2.5
+Chikusei $H$-VNIR-C 343–1018 128 2517×2335 2.5
+Eagle AsiaEAGLE II 401–999 128 2082×1606 1
+Berlin Unknown 455–2447 111 6805×1830 3.6
+Urban Hydice 400-2500 210 307×307 2
+APEX Unknown 350-2500 285 1000×1500 2
+EO-1 Hyperion 357-2567 242 3471×991 30
 Table 5: Properties of 13 Natural Scene and Remote Sensing Hyperspectral Datasets.
@@
 Poisson Denoising. The pre-trained model was directly tested on this task with a Poisson noise intensity scaling factor of 10.
-[…]
-[…]
-[…]
-[…]
+Type Methods Gaussian Denoising (ICVL [3]) Gaussian Denoising (ARAD [4]) Gaussian Denoising (Xiong’an [72])
+Sigma = 30 Sigma = 50 Sigma = 70 Sigma = 30 Sigma = 50 Sigma = 70 Sigma = 30 Sigma = 50 Sigma = 70
+PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$
+Task Specific QRNN3D [63] 42.18 / 0.967 39.70 / 0.942 38.09 / 0.933 41.67 / 0.967 39.15 / 0.935 36.71 / 0.894 37.86 / 0.870 36.03 / 0.825 34.29 / 0.792
 SST [33] 43.32 / 0.976 41.09 / 0.952 39.51 / 0.949 43.02 / 0.972 40.58 / 0.951 38.99 / 0.941 39.26 / 0.878 37.34 / 0.848 35.99 / 0.824
 SERT [34] 43.53 / 0.978 41.32 / 0.966 39.82 / 0.956 43.21 / 0.975 40.84 / 0.959 39.21 / 0.945 39.54 / 0.885 37.58 / 0.859 36.37 / 0.833
-[…]
-[…]
+LDERT [35] 44.12 / 0.982 41.68 / 0.968 39.95 / 0.957 43.74 / 0.979 41.35 / 0.966 39.32 / 0.950 39.92 / 0.889 37.96 / 0.868 36.54 / 0.838
+All in One AirNet [32] 42.02 / 0.966 39.68 / 0.942 37.59 / 0.923 41.39 / 0.963 39.08 / 0.933 37.09 / 0.903 34.04 / 0.700 31.61 / 0.665 30.17 / 0.639
 PromptIR [52] 42.40 / 0.971 40.14 / 0.954 38.20 / 0.934 41.84 / 0.967 39.55 / 0.947 37.67 / 0.921 34.90 / 0.715 32.76 / 0.680 31.31 / 0.657
@@
 PromptHSI [30] 42.61 / 0.976 40.27 / 0.960 39.08 / 0.945 41.90 / 0.971 39.84 / 0.959 38.37 / 0.938 39.54 / 0.902 37.80 / 0.877 36.87 / 0.864
-[…]
+MP-HSIR (Ours) 43.62 / 0.977 41.41 / 0.963 39.82 / 0.951 43.12 / 0.975 40.88 / 0.960 39.28 / 0.946 40.55 / 0.922 38.70 / 0.896 37.17 / 0.874
 Table 6: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different Gaussian noise levels on Gaussian denoising tasks.
 The best and second-best performances are highlighted in red and blue, respectively.
-[…]
-[…]
-[…]
-[…]
+Type Methods Complex Denoising (ICVL [3]) Complex Denoising (ARAD [4]) Complex Denoising (WDC [81])
+Case = 1 Case = 2 Case = 3 Case = 4 Case = 1 Case = 2 Case = 3 Case = 4 Case = 1 Case = 2 Case = 3 Case = 4
+PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$
+Task Specific QRNN3D [63] 42.24 / 0.969 41.98 / 0.968 41.62 / 0.968 40.55 / 0.960 41.74 / 0.966 41.55 / 0.965 41.38 / 0.964 39.76 / 0.946 31.98 / 0.885 31.77 / 0.882 31.48 / 0.878 28.05 / 0.822
 SST [33] 43.38 / 0.976 42.69 / 0.973 42.51 / 0.972 41.16 / 0.964 42.84 / 0.973 42.38 / 0.971 42.01 / 0.970 40.56 / 0.956 33.85 / 0.907 33.69 / 0.905 33.37 / 0.901 29.92 / 0.841
 SERT [34] 43.96 / 0.980 43.48 / 0.978 43.45 / 0.977 42.37 / 0.969 43.56 / 0.978 43.19 / 0.976 42.88 / 0.974 41.85 / 0.963 34.48 / 0.922 34.26 / 0.920 33.98 / 0.915 30.53 / 0.854
-[…]
-[…]
+LDERT [35] 44.04 / 0.981 43.57 / 0.979 43.55 / 0.979 42.51 / 0.971 43.67 / 0.979 43.33 / 0.977 43.06 / 0.974 42.02 / 0.965 34.65 / 0.923 34.42 / 0.920 34.13 / 0.917 30.74 / 0.856
+All in One AirNet [32] 42.11 / 0.968 41.24 / 0.964 40.89 / 0.961 38.49 / 0.942 41.62 / 0.965 40.83 / 0.959 40.31 / 0.957 37.59 / 0.905 29.02 / 0.752 28.93 / 0.744 28.67 / 0.740 25.67 / 0.667
 PromptIR [52] 42.76 / 0.973 41.93 / 0.969 41.43 / 0.969 39.04 / 0.947 42.26 / 0.969 41.54 / 0.964 40.90 / 0.964 38.12 / 0.919 29.84 / 0.761 29.71 / 0.754 29.39 / 0.747 26.38 / 0.676
@@
 PromptHSI [30] 40.61 / 0.967 40.36 / 0.965 39.30 / 0.960 36.27 / 0.927 40.22 / 0.951 39.98 / 0.955 39.42 / 0.953 35.39 / 0.884 34.93 / 0.931 34.87 / 0.930 34.49 / 0.928 30.78 / 0.857
-[…]
+MP-HSIR (Ours) 43.07 / 0.975 42.46 / 0.972 42.20 / 0.972 41.42 / 0.966 42.74 / 0.973 42.32 / 0.971 41.93 / 0.970 40.98 / 0.960 35.21 / 0.933 34.99 / 0.931 34.72 / 0.928 31.36 / 0.880
 Table 7: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different cases on Complex denoising tasks.
@@
 As illustrated in the figure, both PromptIR and InstructIR struggle to effectively differentiate between various HSI degradation types, whereas our method demonstrates distinct separation for each individual task, highlighting its superior capability in handling diverse degradation scenarios.
-[…]
-[…]
-[…]
-[…]
+Type Methods Gaussian Deblurring (ICVL [3]) Gaussian Deblurring (PaviaC [24]) Gaussian Deblurring (Eagle [51])
+Radius = 9 Radius = 15 Radius = 21 Radius = 7 Radius = 11 Radius = 15 Radius = 7 Radius = 11 Radius = 15
+PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$
+Task Specific Stripformer [58] 49.21 / 0.994 45.85 / 0.989 43.04 / 0.980 38.73 / 0.934 36.95 / 0.914 35.72 / 0.891 44.68 / 0.974 41.37 / 0.958 39.83 / 0.942
 FFTformer [27] 49.83 / 0.994 46.43 / 0.989 43.69 / 0.981 39.82 / 0.942 37.63 / 0.919 36.44 / 0.901 45.51 / 0.980 42.18 / 0.962 40.59 / 0.944
 LoFormer [47] 50.35 / 0.995 46.94 / 0.991 44.15 / 0.983 39.54 / 0.941 37.41 / 0.918 36.21 / 0.892 45.32 / 0.979 42.03 / 0.961 40.42 / 0.946
-[…]
-[…]
+MLWNet [14] 50.68 / 0.996 47.54 / 0.991 44.76 / 0.983 40.85 / 0.949 38.92 / 0.929 37.28 / 0.901 46.61 / 0.983 43.68 / 0.969 41.83 / 0.955
+All in One AirNet [32] 50.64 / 0.995 47.10 / 0.991 43.89 / 0.982 39.45 / 0.940 37.76 / 0.921 36.08 / 0.893 44.83 / 0.977 42.32 / 0.966 40.11 / 0.946
 PromptIR [52] 51.00 / 0.996 47.70 / 0.992 44.32 / 0.983 40.41 / 0.948 38.62 / 0.929 37.13 / 0.904 46.17 / 0.981 43.45 / 0.968 41.60 / 0.955
@@
 PromptHSI [30] 25.16 / 0.619 30.89 / 0.852 41.52 / 0.980 38.73 / 0.936 36.63 / 0.913 34.84 / 0.888 42.45 / 0.972 39.34 / 0.952 37.46 / 0.938
-[…]
+MP-HSIR (Ours) 51.53 / 0.996 47.60 / 0.992 45.07 / 0.982 40.85 / 0.949 38.95 / 0.928 37.19 / 0.905 46.26 / 0.982 43.54 / 0.969 41.36 / 0.952
 Table 8: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different blur kernel radius on Gaussian deblurring tasks.
@@
 In this section, we present additional experimental results, including the results of more quantitative results, model efficiency, more results of ablation study, controllability analysis, and more visual comparisons.
-[…]
-[…]
-[…]
-[…]
+Type Methods Super-Resolution (ARAD [4]) Super-Resolution (PaviaU [24]) Super-Resolution (Houston [66])
+Scale = 2 Scale = 4 Scale = 8 Scale = 2 Scale = 4 Scale = 8 Scale = 2 Scale = 4 Scale = 8
+PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$
+Task Specific SNLSR [23] 43.93 / 0.980 34.56 / 0.902 29.67 / 0.813 34.58 / 0.869 29.85 / 0.719 27.23 / 0.601 34.91 / 0.908 31.33 / 0.782 28.86 / 0.671
 MAN [60] 44.81 / 0.985 35.35 / 0.912 30.49 / 0.830 34.92 / 0.872 30.26 / 0.723 27.59 / 0.604 35.26 / 0.911 31.68 / 0.785 29.15 / 0.677
 ESSAformer [78] 45.32 / 0.988 36.02 / 0.927 30.85 / 0.838 35.47 / 0.879 30.60 / 0.728 27.96 / 0.606 35.60 / 0.913 31.94 / 0.787 29.57 / 0.679
-[…]
-[…]
+SRFormer [80] 45.84 / 0.989 36.73 / 0.931 31.48 / 0.845 35.92 / 0.887 31.08 / 0.745 28.41 / 0.620 36.15 / 0.920 32.44 / 0.805 29.81 / 0.684
+All in One AirNet [32] 44.82 / 0.985 35.26 / 0.919 30.12 / 0.828 34.85 / 0.871 30.19 / 0.724 27.76 / 0.609 35.22 / 0.911 31.65 / 0.786 29.13 / 0.676
 PromptIR [52] 45.33 / 0.988 36.00 / 0.927 30.77 / 0.838 35.57 / 0.883 30.81 / 0.735 28.20 / 0.619 35.96 / 0.915 32.38 / 0.798 29.85 / 0.684
@@
 PromptHSI [30] 40.25 / 0.975 35.41 / 0.931 29.35 / 0.806 34.84 / 0.871 30.13 / 0.722 27.27 / 0.602 35.34 / 0.912 31.62 / 0.778 28.59 / 0.635
-[…]
+MP-HSIR (Ours) 46.72 / 0.991 36.88 / 0.939 31.14 / 0.843 36.27 / 0.894 31.26 / 0.757 28.38 / 0.630 36.57 / 0.926 32.68 / 0.813 29.92 / 0.690
 Table 9: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different downsampling scales on Super-Resolution tasks.
@@
 Table 10: No-reference quality assessment on real datasets.
-[…]
-[…]
-[…]
-[…]
+Type Methods Inpainting (ICVL [3]) Inpainting (Chikusei [73])
+Rate = 0.7 Rate = 0.8 Rate = 0.9 Rate = 0.7 Rate = 0.8 Rate = 0.9
+PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$
+Task Specific NAFNet [8] 45.03 / 0.989 44.65 / 0.988 43.50 / 0.985 40.34 / 0.952 40.15 / 0.955 38.97 / 0.952
 Restormer [76] 46.51 / 0.991 46.00 / 0.991 44.85 / 0.988 36.52 / 0.902 36.34 / 0.892 36.13 / 0.903
 DDS2M [48] 45.41 / 0.989 43.34 / 0.983 37.80 / 0.935 36.77 / 0.906 35.23 / 0.901 32.83 / 0.854
-[…]
-[…]
+HIR-Diff [50] 41.82 / 0.973 38.87 / 0.949 36.04 / 0.924 38.59 / 0.923 37.96 / 0.920 36.41 / 0.904
+All in One AirNet [32] 43.10 / 0.983 43.06 / 0.983 41.65 / 0.977 38.12 / 0.919 37.86 / 0.921 36.39 / 0.918
 PromptIR [52] 46.96 / 0.992 46.93 / 0.992 45.24 / 0.988 38.86 / 0.925 38.30 / 0.931 37.05 / 0.934
@@
 PromptHSI [30] 42.83 / 0.983 41.72 / 0.976 39.89 / 0.956 38.99 / 0.966 37.64 / 0.952 35.35 / 0.920
-[…]
+MP-HSIR (Ours) 53.06 / 0.997 51.94 / 0.996 49.60 / 0.994 44.75 / 0.981 44.06 / 0.981 42.08 / 0.975
 Table 11: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different mask rates on Inpainting tasks.
 The best and second-best performances are highlighted in red and blue, respectively.
-[…]
-[…]
-[…]
-[…]
+Type Methods Dehazing (PaviaU [24]) Dehazing (Eagle [51])
+Omega = 0.5 Omega = 0.75 Omega = 1.0 Omega = 0.5 Omega = 0.75 Omega = 1.0
+PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$
+Task Specific SGNet [45] 36.52 / 0.974 34.10 / 0.964 32.22 / 0.949 39.43 / 0.989 37.33 / 0.976 34.90 / 0.962
 SCANet [18] 39.01 / 0.986 36.54 / 0.978 34.21 / 0.969 41.92 / 0.991 39.68 / 0.987 37.31 / 0.978
 MB-Taylor [54] 40.51 / 0.991 38.03 / 0.984 35.44 / 0.975 43.76 / 0.995 40.97 / 0.992 38.36 / 0.986
-[…]
-[…]
+DCMPNet [79] 39.63 / 0.993 37.14 / 0.985 34.82 / 0.976 42.93 / 0.995 40.15 / 0.991 37.64 / 0.985
+All in One AirNet [32] 38.61 / 0.982 35.65 / 0.967 32.51 / 0.947 41.88 / 0.991 38.68 / 0.982 35.92 / 0.969
 PromptIR [52] 40.34 / 0.991 37.43 / 0.983 34.47 / 0.971 43.55 / 0.995 40.69 / 0.992 37.94 / 0.988
@@
 PromptHSI [30] 38.62 / 0.982 36.48 / 0.975 35.22 / 0.964 40.88 / 0.986 40.49 / 0.984 37.98 / 0.981
-[…]
+MP-HSIR (Ours) 42.64 / 0.993 39.46 / 0.988 36.68 / 0.978 45.66 / 0.997 42.24 / 0.995 39.34 / 0.992
 Table 12: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different haze levels on Dehazing tasks.
 The best and second-best performances are highlighted in red and blue, respectively.
-[…]
-[…]
-[…]
-[…]
+Type Methods Band Completion (ARAD [4]) Band Completion (Berlin [49])
+Rate = 0.1 Rate = 0.2 Rate = 0.3 Rate = 0.1 Rate = 0.2 Rate = 0.3
+PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$ PSNR / SSIM $\uparrow$
+Task Specific NAFNet [8] 47.82 / 0.996 47.02 / 0.995 46.29 / 0.994 39.71 / 0.972 37.51 / 0.874 37.84 / 0.875
 Restormer [76] 49.24 / 0.997 48.29 / 0.995 47.49 / 0.993 34.98 / 0.605 35.24 / 0.606 35.00 / 0.607
 SwinIR [37] 50.95 / 0.997 49.80 / 0.995 48.49 / 0.993 36.88 / 0.950 34.69 / 0.855 34.78 / 0.854
-[…]
-[…]
+MambaIR [16] 51.46 / 0.998 50.32 / 0.995 49.01 / 0.993 37.54 / 0.953 35.38 / 0.857 35.43 / 0.855
+All in One AirNet [32] 46.14 / 0.994 45.21 / 0.992 44.46 / 0.990 37.62 / 0.691 35.26 / 0.586 34.86 / 0.595
 PromptIR [52] 47.68 / 0.996 46.71 / 0.994 45.41 / 0.992 42.81 / 0.707 39.55 / 0.910 39.00 / 0.640
@@
 PromptHSI [30] 49.05 / 0.996 47.09 / 0.993 45.89 / 0.992 47.11 / 0.997 43.14 / 0.973 39.82 / 0.956
-[…]
+MP-HSIR (Ours) 57.83 / 0.999 56.61 / 0.999 54.99 / 0.998 52.14 / 0.999 49.20 / 0.997 47.26 / 0.965
 Table 13: [All-in-one] Quantitative comparison of all-in-one and state-of-the-art task-specific methods under different mask rates on Band Completion tasks.
@@
 The results on the Urban and EO-1 datasets are reported in Table 10, providing a more comprehensive assessment of generalization ability under practical conditions.
-[…]
-[…]
-[…]
+Methods Natural Scene Remote Sensing
+Params (M) FLOPS (G) Params (M) FLOPS (G)
+AirNet [32] 5.82 19.04 12.23 43.79
 PromptIR [52] 33.00 10.03 72.60 22.21
@@
 PromptHSI [30] 25.90 10.10 50.89 21.91
-[…]
+MP-HSIR (Ours) 13.88 14.40 30.91 32.74
 Table 14: Model complexity comparisons